Horizon Value of Free Cash Flows

Current and projected free cash flows for Radell Global Operations are shown below.

Actual
2016

2017
Projected
2018

2019
Free cash flow $606.82 $667.50 $707.55 $750.00
(millions of dollars)

Growth is expected to be constant after 2018, and the weighted average cost of capital is 11.5%. What is the horizon (continuing) value at 2019 if growth from 2018 remains constant? Round your answer to the nearest dollar. Do not round intermediate calculations.

$_________?
Answer #1

Growth =(750-707.55)/707.55 =6%
Horizon Value =FCF 2019*(1+g)/(Cost of capital-growth) =750*(1+6%)/(11.5%-6%) =14454.55